OpenStudy (anonymous):

Find equations of the lines passing through (-2,4) and having the following characteristics. a) Slope of 7/16 b) parallel to the line 5x-3y=3 c) Passing through the origin. d) Parallel to the y-axis

OpenStudy (anonymous):

a) You know y= mx + b, you are given the slope.. m= 7/16 so now just plug in a know point with x and y values (-2,4) to find b (the point where the line crosses the y-axis) y = (7/16)x + b 4 = (7/16) (-2) + b Solve for b, and the equation of your line should be y= (7/16)x + (insert value of b here)

OpenStudy (anonymous):

Oh kk thanx

OpenStudy (anonymous):

Do you need help on the other ones?

OpenStudy (anonymous):

No i think i get them now thank you anyway.